OpenStudy (anonymous):

How many grams of magnesium metal will react completely with 5.2 liters of 4.0 M HCl? Mg (s) + 2 HCl (aq) --> MgCl2 (aq) + H2 (g)

OpenStudy (anonymous):

Since we know that there are 2 moles of HCl for every one mole of Mg metal all we have to do is find the moles of HCl then use the stoichiometric ratio then convert moles to grams. \[4.0M \times 5.2L = 20.8moles\]\[20.8molesHcl \times {1moleMg \over 2molesHCl } = 10.4moles Mg\]\[10.4molesMg \times {24.305grams \over moles} = 2.5 x10^{1} grams\]